# Leveraging Large Language Models for Collective Decision-Making

### v2

@@ Acknowledgements.
We want to thank the researchers at Microsoft’s Office of Applied Research – and, in particular, Bahar Sarrafzadeh, Victor Poznanski, Yujin Kim, and Aaron Halfaker – for the stimulating discussions and feedback on our project. We would also like to thank Shilad Sen for providing us with the synthetic dataset for meeting scheduling.
@@ A.1. Member Data
Name
Manager
Role
Level
Responsibilities
Schedule Preferences
Member 1
CEO
5
[’Setting the vision, mission, and values of the company’, ”Leading the overall strategy and execution of the company’s initiatives”, ’Managing the financial, legal, and operational aspects of the company’]
She prefers to have meetings in the morning and block out time for deep work in the afternoon. She is flexible with her availability, but appreciates advance notice and clear agendas. She uses a shared calendar to keep track of her appointments and tasks.
Member 34
Member 1
Senior Customer Success Manager
3
[”Managing a portfolio of key accounts and ensuring they achieve their desired outcomes and satisfaction with the company’s products and services”, ’Providing proactive and reactive support, guidance, and training to customers’, ’Identifying and pursuing opportunities to upsell, cross-sell, and renew customers’]
He likes to have meetings throughout the day and adjust his schedule according to his customers’ preferences and availability. He is flexible with his availability, but prefers to have meetings scheduled at least a day in advance. He uses a customer success platform to manage his customer interactions and outcomes.
Member 33
Member 1
Senior Software Engineer
3
[”Designing, developing, and testing software solutions for the company’s products and services”, ’Troubleshooting and resolving software issues and bugs’, ’Mentoring and coaching junior software engineers and sharing best practices’]
He prefers to have meetings in the morning and use the rest of the day for coding and debugging. He is flexible with his availability, but prefers to have meetings scheduled at least a few hours in advance. He uses a code review tool to collaborate and communicate with his peers and managers.
Member 29
Member 22
Sales Representative
2
[”Identifying and qualifying leads and prospects for the company’s products and services”, ”Conducting sales presentations and demos to potential customers and explaining the value proposition and benefits of the company’s products and services”, ’Following up and closing sales deals and ensuring customer satisfaction and retention’]
She likes to have meetings in the afternoon and evening, and use the morning and early afternoon for prospecting and following up with leads and customers. She is flexible with her availability, but likes to have meetings scheduled at least a day in advance. She uses a CRM tool to manage her sales pipeline and activities.
@@ A.2. Meeting Data
Organizer
Members
Subject Date Duration
Member 34
Member 1, Member 29, Member 33
Discuss Customer Success Efforts 2/16/2023 30 minutes
Member 14
Member 13, Member 33, Member 15, Member 4, Member 11
AI feature update and feedback 2/15/2023 30 minutes
@@ B.1. Intent Extraction Prompts
The first prompt extracts the members’ messages $m_{i}^{t}$ from a member $i$ at round $t$. The prompt is given below:
@@ B.2. Coordination Module Prompt
We present the coordination module prompt for rounds $2\leq t\leq T$ which corresponds to the most general setting:
@@ Appendix C Simulation Plots
Figure 10. Results from simulations from the study where the schedule preferences and the knowledge graph are taken into account.
@@ D.4. Synthetic Meeting Scenario
Survey Questions. Below, we give the questions and the potential answer choices (if they exist) used in our survey, for each of the three stages of the survey:

### v3

@@ Acknowledgements.
We want to thank the researchers at Microsoft’s Office of Applied Research – particularly Bahar Sarrafzadeh, Victor Poznanski, Yujin Kim, Katrina Zuccaro, and Aaron Halfaker – for stimulating discussions and feedback on our project. We would also like to thank Shilad Sen for providing us with the synthetic dataset for meeting scheduling. Finally, M.P. would like to thank Jon Kleinberg and Farhana Shahid for discussing the early versions of the manuscript.
@@ A.1. Member Data
Name Manager Role Level Responsibilities Schedule Preferences
Member 1 CEO 5 [’Setting the vision, mission, and values of the company’, ”Leading the overall strategy and execution of the company’s initiatives”, ’Managing the financial, legal, and operational aspects of the company’] She prefers to have meetings in the morning and block out time for deep work in the afternoon. She is flexible with her availability but appreciates advance notice and clear agendas. She uses a shared calendar to keep track of her appointments and tasks.
Member 34 Member 1 Senior Customer Success Manager 3 [”Managing a portfolio of key accounts and ensuring they achieve their desired outcomes and satisfaction with the company’s products and services”, ’Providing proactive and reactive support, guidance, and training to customers’, ’Identifying and pursuing opportunities to upsell, cross-sell, and renew customers’] He likes to have meetings throughout the day and adjust his schedule according to his customers’ preferences and availability. He is flexible with his availability, but prefers to have meetings scheduled at least a day in advance. He uses a customer success platform to manage his customer interactions and outcomes.
Member 33 Member 1 Senior Software Engineer 3 [”Designing, developing, and testing software solutions for the company’s products and services”, ’Troubleshooting and resolving software issues and bugs’, ’Mentoring and coaching junior software engineers and sharing best practices’] He prefers to have meetings in the morning and use the rest of the day for coding and debugging. He is flexible with his availability, but prefers to have meetings scheduled at least a few hours in advance. He uses a code review tool to collaborate and communicate with his peers and managers.
Member 29 Member 22 Sales Representative 2 [”Identifying and qualifying leads and prospects for the company’s products and services”, ”Conducting sales presentations and demos to potential customers and explaining the value proposition and benefits of the company’s products and services”, ’Following up and closing sales deals and ensuring customer satisfaction and retention’] She likes to have meetings in the afternoon and evening, and use the morning and early afternoon for prospecting and following up with leads and customers. She is flexible with her availability, but likes to have meetings scheduled at least a day in advance. She uses a CRM tool to manage her sales pipeline and activities.
@@ A.2. Meeting Data
Organizer Members Subject Date Duration
Member 34 Member 1, Member 29, Member 33 Discuss Customer Success Efforts 2/16/2023 30 minutes
Member 14 Member 13, Member 33, Member 15, Member 4, Member 11 AI feature update and feedback 2/15/2023 30 minutes
@@ B.1. Intent Extraction Prompts
The first prompt instructs the LLM how to communicate with member $i$ at any round $t$. The prompt is given below:
@@ B.2. Coordination Module Prompt
We present the coordination module prompt for rounds $2\leq t\leq T$, which corresponds to the most general setting:
@@ Appendix C Simulation Plots
Figure 10. Results from simulations from the study where the schedule preferences and the database are taken into account.
@@ D.4. Synthetic Meeting Scenario
Survey Questions. Below, we give the questions and the potential answer choices (if they exist) used in our survey for each of the three stages of the survey:
@@
Appendix E Administrator Interview Questions
  1. (1)
    What types of events do you typically coordinate (e.g., meetings, team outings)? Which occurs most frequently?
  2. (2)
    On average, how much time per week do you spend coordinating events (e.g., emails, gathering responses, scheduling)? What percentage of your workweek is this?
  3. (3)
    How many events do you typically coordinate in a week? (Include any events you are responsible for, even if ongoing.)
  4. (4)
    How many participants do you typically communicate with per event?
  5. (5)
    What percentage of participants are from external organizations on average?
  6. (6)
    Do you find any challenges when coordinating with external participants? Please briefly share the difference between and coordinating completely within your organization.
  7. (7)
    What tools do you use for event coordination? Briefly describe your experience, listing the pros and cons.
  8. (8)
    What factors do you consider when coordinating events (e.g., power dynamics, location)? Please list them.
  9. (9)
    How do you balance these factors? What is your decision-making process?
  10. (10)
    After gathering participants’ input, how many options do you typically suggest?
  11. (11)
    How do you handle situations where no option works for all participants? Briefly describe your approach.
  12. (12)
    On average, how many times do you propose new options for an event?
  13. (13)
    How often does the final plan not work for everyone?
  14. (14)
    What percentage of the initial participants typically agree to or attend the event?
  15. (15)
    How confident are you that participants’ preferences are considered equitably when coordinating events?
  16. (16)
    What challenges or frustrations do you face during event coordination? What assistance would improve this process?